# Comparing Causal Inference Methods for Point Exposures with Missing Confounders: A Simulation Study

### v3

@@ 2.1 Setting and Notation
Let $R\equiv(R_{1},\ldots,R_{q})\in\{0,1\}^{q}$ be a vector of binary indicators denoting whether or not each of the $q$ partially observed confounders $L_{p}$ is observed (i.e., $R_{j}=1$ if and only if the $j$-th component of $L_{p}$ is observed), with $L_{p}^{(r)}$ representing the components of $L_{p}$ that are actually observed under the observation pattern $R=r$. As in Levis et al.[1], we take the observed data to be $n$ i.i.d. observations of $O=(L_{c},A,Y,R,L_{p}^{(R)})\sim P_{o}$. Defining $S=\mathds{1}(R=\bm{1}_{q})$ to be a complete data indicator, we can also consider the coarsened observed data, $O^{\prime}=(L_{c},A,Y,S,SL_{p})\sim P_{o^{\prime}}$. When $S=0$, we conceive of $SL_{p}=0$ to denote that the entire $L_{p}$ is treated as missing (e.g., NA) Note that both $P_{o}$ and $P_{o^{\prime}}$ are determined jointly by $P_{f}$, the distribution of the full data $(L,A,Y)$, and the missingness mechanism that underpins $R$ [1].
@@ 2.3 Estimators Based on a Complete-Case MAR Assumption
To paraphrase Levis et al.[1]: $\eta(L_{c},a)=P_{o^{\prime}}[A=a~{}|~{}L_{c}]$ is a model for treatment mechanism conditional on only the subset of confounders that are observed for all subjects; $\mu(y~{}|~{}L_{c},A)=P_{o^{\prime}}[Y\leq y~{}|~{}L_{c},A]$ is a outcome model based only on the treatment and the subset of confounders that are observed for all subjects; $\pi(L_{c},A,Y)=P_{o^{\prime}}[S=1~{}|~{}L_{c},A,Y]$ is a model for the probability of being a “complete case”, which only depends on treatment, outcome, and complete confounders – quantities that are observable for all subjects; and finally $\lambda(\ell_{p}~{}|~{}L_{c},A,Y,S=1)$ is the conditional density of $L_{p}$ given the remaining observed variables, among complete cases for which no confounders are missing. The $S$ in the exponent above the $\lambda$ term in Equation (1) sets that portion of the coarsened observed data likelihood to $1$ if $L_{p}$ are not observed. Table 2 reproduces Table 1 from Levis et al.[1], summarizing these nuisance functions. A visual overview of this novel data factorization is given in Figure 2(a).
@@
On the basis of this novel factorization of the observed (coarsened) data likelihood, Levis et al.[1] introduce a complete-case missing at random (CCMAR) assumption, and propose two estimators for the mean counterfactual $\mathbb{E}[Y(a)]$ that are valid under CCMAR. The first is an inverse-weighted outcome regression (IWOR) estimator, $\tilde{\chi}_{a}$, and the second is an influence function (IF) based estimator $\hat{\chi}_{a}$. The influence function-based estimator, $\hat{\chi}_{a}$, is of primary interest as Levis et al.[1] are able to prove several results regarding the robustness and non-parametric efficiency of this estimator. Nevertheless, $\tilde{\chi}_{a}$ serves as a useful comparison point in our simulation study as another estimator derived from the factorization considered in Equation (1). Formal definitions and derivations of $\tilde{\chi}_{a}$ and $\hat{\chi}_{a}$ are presented in §$\S 5.1$ and §$\S 5.2$ of Levis et al.[1], respectively; for completeness, we give the form of these estimators in the supplementary materials. Henceforth, we will refer to $\tilde{\chi}_{a}$ and $\hat{\chi}_{a}$ as CCMAR-based estimators.
@@
A noteworthy feature of this particular assumption is that it is agnostic to which components of $L_{p}$ may be missing for any particular subject. That is, this assumption is strictly in regards to whether a subject has complete data $(S=1)$ or not $(S=0)$. Of immediate consequence is that when this assumption holds, estimation strategies for $\tilde{\chi}_{a}$ and $\hat{\chi}_{a}$ do not require estimation or specification of complex, non-monotone missingness patterns that are common in EHR data[7]. Possible limitations of this coarsening are discussed in detail in §$\S 7$ of Levis et al.[1]. A full set of identification assumptions for the ATE in this setting is covered in §$\S 3.1$ and §$\S 3.2$ of [1] and provided in our Supplementary Materials as well.
@@ 3 Data Driven Simulation Study
    A secondary goal of this simulation study is to evaluate the CCMAR-based estimators perform when data are generated by the alternative factorization in Equation (3). In such settings, the performance of CCMAR-based estimators is more closely aligned with how such estimators are likely to fair in real world settings where component nuisance models are unknown to the analyst.
@@ 3.2 Methods
We considered three imputation models for $L_{p}$: the correctly specified DGP for $L_{p}$ (a gamma generalized linear model for comorbidity score $(L^{(4)})$ and a logistic regression for smoking status $(L^{(5)})$, with exact models depending on the corresponding factorization under which data were generated); a simpler imputation model with a Gaussian linear model for comorbidity score and a logistic regression for smoking status, with neither model considering any interactions; and, an imputation model with a Gaussian linear model for comorbidity score and a logistic regression for smoking status, while specifying all pairwise interactions. To avoid overfitting/variance inflation due to specifying all pairwise interactions, linear/logistic regression models in this third case were fit using glmnet[32] in R with $L_{1}$ (LASSO)[33] regularization. Additionally, we considered a fourth strategy where the problem of missing data was ignored and subjects with missing these data were dropped, in order to mimic a complete case analysis.
@@ 4.1 Data Generation Process
$A$ $\sim\text{Bernoulli}(0.5)$
$Y~{}|~{}A$$\sim\begin{cases}\text{Beta}(2,4)&\text{if}~{}A=0\\ \text{Beta}(4,2)&\text{if}~{}A=1\end{cases}$
$\text{logit}(P[S=1~{}|~{}A,Y])$ $=-0.35+0.5A+0.18Y+0.05A\times Y$
$\text{logit}(P[L_{p1}=1~{}|~{}A,Y,S=1])$ $=-0.6+0.5A+0.25Y+0.1A\times Y$
$L_{p2}~{}|~{}A,Y,S=1,L_{p1}$ $\sim\mathcal{N}(A+Y+2.5L_{p1}\times Y,1.25^{2})$
@@ 5 Results
    • Most methods have relative uncertainty $<1$ (better than influence function CCMAR-based estimator).
@@
Table 7: Data Driven Simulation Scenario #$\#3$: Factorization in Equation (1), 2 partially missing confounders, amplified interactions between $A,L_{p}$, and $Y$ in joint imputation model, induced skew in distribution of imputation model.

### v4

@@ 2.1 Setting and Notation
[…]
@@ 2.3 Estimators Based on a Complete-Case MAR Assumption
To paraphrase Levis et al.[1]: $\eta(L_{c},a)=P_{o^{\prime}}[A=a~|~L_{c}]$ is a model for treatment mechanism conditional on only the subset of confounders that are observed for all subjects; $\mu(y~|~L_{c},A)=P_{o^{\prime}}[Y\leq y~|~L_{c},A]$ is a outcome model based only on the treatment and the subset of confounders that are observed for all subjects; $\pi(L_{c},A,Y)=P_{o^{\prime}}[S=1~|~L_{c},A,Y]$ is a model for the probability of being a “complete case”, which only depends on treatment, outcome, and complete confounders – quantities that are observable for all subjects; and finally $\lambda(\ell_{p}~|~L_{c},A,Y,S=1)$ is the conditional density of $L_{p}$ given the remaining observed variables, among complete cases for which no confounders are missing. The $S$ in the exponent above the $\lambda$ term in Equation (1) sets that portion of the coarsened observed data likelihood to 111 if $L_{p}$ are not observed. Table 2 reproduces Table 1 from Levis et al.[1], summarizing these nuisance functions. A visual overview of this novel data factorization is given in Figure 2(a).
@@
On the basis of this novel factorization of the observed (coarsened) data likelihood, Levis et al.[1] introduce a complete-case missing at random (CCMAR) assumption, and propose two estimators for the mean counterfactual $\mathbb{E}[Y(a)]$ that are valid under CCMAR. The first is an inverse-weighted outcome regression (IWOR) estimator, $\tilde{\chi}_{a}$, and the second is an influence function (IF) based estimator $\hat{\chi}_{a}$. The influence function-based estimator, $\hat{\chi}_{a}$, is of primary interest as Levis et al.[1] are able to prove several results regarding the robustness and non-parametric efficiency of this estimator. Nevertheless, $\tilde{\chi}_{a}$ serves as a useful comparison point in our simulation study as another estimator derived from the factorization considered in Equation (1). Formal definitions and derivations of $\tilde{\chi}_{a}$ and $\hat{\chi}_{a}$ are presented in §5.1$\S$ 5.1§ 5.1 and §5.2$\S$ 5.2§ 5.2 of Levis et al.[1], respectively; for completeness, we give the form of these estimators in the supplementary materials. Henceforth, we will refer to $\tilde{\chi}_{a}$ and $\hat{\chi}_{a}$ as CCMAR-based estimators.
@@
[…]
@@ 3 Data Driven Simulation Study
    A secondary goal of this simulation study is to evaluate the CCMAR-based estimators perform when data are generated by the alternative factorization in Equation (3). In such settings, the performance of CCMAR-based estimators is more closely aligned with how such estimators are likely to fare in real world settings where component nuisance models are unknown to the analyst.
@@ 3.2 Methods
Under each imputation strategy, a parametric model was estimated for $L_{p}$, and subsequently a single draw was sampled from the induced conditional distribution for each missing observation. The choice to use single imputation rather than multiple imputation helped ease computational given the large number of simulation scenarios and comparator methods considered. Three imputation models for $L_{p}$ were considered: the correctly specified DGP for $L_{p}$ (a gamma generalized linear model for comorbidity score $(L^{(4)})$ and a logistic regression for smoking status $(L^{(5)})$, with exact models depending on the corresponding factorization under which data were generated); a simpler imputation model with a Gaussian linear model for comorbidity score and a logistic regression for smoking status, with neither model considering any interactions; and, an imputation model with a Gaussian linear model for comorbidity score and a logistic regression for smoking status, while specifying all pairwise interactions. To avoid overfitting/variance inflation due to specifying all pairwise interactions, linear/logistic regression models in this third case were fit using glmnet[32] in R with $L_{1}$ (LASSO)[33] regularization. Additionally, we considered a fourth strategy where the problem of missing data was ignored and subjects with missing these data were dropped, in order to mimic a complete case analysis.
@@ 4.1 Data Generation Process
[…]
@@ 5 Results
    • Most methods have relative uncertainty <1<1< 1 (better than influence function CCMAR-based estimator).
@@
Table 7: Data Driven Simulation Scenario #3\#3# 3: Factorization in Equation (1), 2 partially missing confounders, amplified interactions between $A,L_{p}$, and $Y$ in joint imputation model, induced skew in distribution of imputation model.
@@ 6 Discussion
Understanding that no simulation study can hope to cover all possible scenarios, we believe that those that we do consider collectively provide useful insight on key decisions that analysts have to engage in. Furthermore, we believe that the framework we present provides a path for researchers designing their own related simulations, perhaps tailored to settings they find themselves in. Our code is made publicly available at https://github.com/lbenz730/missing_confounder_sims/, and we hope researchers interested in additional scenarios or other alternatives can use our resources as a starting point to run their own simulations.
@@
Though we used single imputation rather than multiple imputation to lower the computational burden of our simulation study, we anticipate similar conclusions from using multiple imputation. Since our imputation scheme is equivalent to multiple imputation with $M=1$, increasing the number of imputed datasets will not change bias properties of the estimators considered, though the variability of estimators across simulated datasets may be slightly smaller. Because non-trivial bias remained in many scenarios even when missing data were imputed by the true data generating process, more complex imputation methods are not a remedy for misspecification of models necessary to adjust for confounding. Nevertheless, methods like multiple imputation by chained equations (MICE) [45], joint modeling multiple imputation (JOMO) [46], and substantive model compatible fully conditional specification (SMCFCS) [47] may be useful tools for consideration in more complex settings when $L_{p}$ is high dimensional or missingness patterns are non-monotone. Practically, however, state-of-the-art methods for non-monotone MAR data [27] rely on a set of parametric models for the missingness probabilities, which unfortunately involve unobservable strata. Additional efficiency considerations based on this direction are discussed in more detail in Levis et al. [1], but such discussion is outside of the primary scope of our work.